OpenStudy (anonymous):

Find the vertex and the axis of symmetry for the parabola with the given equation. y=5x^2-30x+47

OpenStudy (anonymous):

vertex: (3,2) symmetry x=3

OpenStudy (dumbcow):

Use x = -b/2a

OpenStudy (anonymous):

yes do that

OpenStudy (saifoo.khan):

Right

OpenStudy (anonymous):

x=-b/(2a) for axis of symmetry ax^2+bx+c b=-30 a=5 x=-(-30)/(2*5)=30/10=3 solve for y 5x^2-30x+47 5(3)^2-30(3)+47 5*9+47-90 45+47-90 92-90 y=2